RC meeting and Run Plan, Monday October 24th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today to accommodate Hall C weekly meeting

  • Location: Counting house conference room

Useful Links

Current Status

Vacuum Issues

  • The vacuum techs were concerned about the turbo pump VTP1H00A --- EPICS read back was trending down from 1500 HZ to 1470 HZ now.
  • Jessie visually checked the turbo pump and discovered it did appear to be failing: Decreasing rotation frequency and a large current read back. With the assistance of the vacuum techs to monitor the upstream vacuum, Jessie replaced the turbo pump and controller [1]
  • Image of the vacuum and valves along beam line: [2]
  • SWING Shift (18:30): Returned to Beam Permit and began to check the beam profile and raster.

Polarized Target

  • Target Polarization was reduced after EPR measurement taken during SWING shift: https://logbooks.jlab.org/entry/4066891
  • With the lower polarization, decided to perform a "spin up".
    • How did it go?

Carbon Optics

  • While waiting for the polarization to increase, Simona worked on understanding the high ion chamber readings found during the attempted SBS GEM luminosity scan.
    • The scan couldn't reach above 3 uA without tripping the ion chambers: [3]
  • elog on scrapping with carbon optics target: [4]
  • The rates and ion chamber readings for the carbon hole and the polarized helium-3 cell are the same as they were on October 16: [5]

Other Topics

  • High data rate: Increased the BBCal threshold from 282 to 366 and set ps1=-1
  • Currently there is no planned RV recovery or beam studies for this week
  • OPS and RF would like to study the ramp rate in Hall C while observing their ramp rate. This can be done in a dedicated beam studies, in which "parasitic beam" could be delivered to the halls. Jay anticipates the "parasitic beam" would be similar to what we are receiving now for production but likely with a higher trip rate. The test duration would be no more than 4 hours.
    • If it cannot be done due to Hall A's ramp rate (1 uA/s), they wonder if the ramp on the polarized helium-3 cell could be raised? To ~1.5 uA/s? Switch to a non-glass cell during this period?
  • Excel spreadsheet created to track the DAQ crashes their sources was setup.
  • Subsystem updates?
  • A collaborator needs a hall walk through. Are there any others than may need one in order to coordinate with Jessie.
  • Configuration and pass change moved to Monday (10/31/2022): [6]
    • Swapping of the leads to the power supply of the 3rd Moller quad
  • Opportunistic BCM calibration (higher current) before pass change?

Current Run Plan

For Monday swing and Tuesday OWL we will be focusing on polarized helium-3 production.

Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam delivery is interrupted for more than 4 hours, perform a checkout of the beam profile and raster follow Simona's procedure: [7]
  • Cross check the rate with what is expected: [8]
  • Cross check with the expected ion chamber readings: [9]

Notices in Effect for Shift Crew

Current Issues

  • Shift Openings: [10]


Planned Accesses to the Hall

Special Requests